### v1

# Weakly-supervised diagnosis identification from Italian discharge letters

Abstract
Objective: Identifying the diagnosis that determines a patient’s hospitalization or Emergency Room admission is crucial for cohort selection and epidemiological analyses. Discharge letters are the only type of data consistently produced across hospitals. Recognizing diseases from discharge letters is a classic document classification problem, typically requiring supervised learning. However, manual annotation of large datasets of discharge letters is uncommon since it is extremely time-consuming. We propose a novel weakly-supervised pipeline to recognize diseases from Italian discharge letters.
Methods: Our Natural Language Processing pipeline is based on a fine-tuned version of the Italian Umberto model. The pipeline extracts diagnosis-related sentences from a subset of letters and applies a two-level clustering using the embeddings generated by the fine-tuned Umberto model. These clusters are summarized and those mapped to the diseases of interest are selected as weak labels. Finally, the same BERT-based model is trained using these weak labels to detect the targeted diseases.
Results: Results on a case study related to the identification of bronchiolitis in a dataset of 33’176 Italian discharge letters from multiple hospitals in the Veneto Region demonstrate the potential of our method, with an AUC of 77.7 % and an F1-Score of 75.1 % on manually annotated labels, improving compared to other non-supervised methods and with a limited loss compared to fully supervised methods. Results are robust to the cluster selection and the identified clusters highlight the potential to recognize a variety of diseases.
Conclusions: This study demonstrates the feasibility of successfully addressing the challenge of diagnosis identification from Italian discharge letters in the absence of labelled data. Our weakly-supervised pipeline showed strong performance and robustness, and its flexibility allows for easy adaptation to various diseases. This approach offers a scalable solution for clinical text classification, significantly reducing the need for manual annotation while maintaining good accuracy.
keywords:
Natural Language Processing , weakly-supervised learning , discharge letters , diagnosis identification
MSC:
68T50
[…]
[inst1]organization=MOX - Modelling and Scientific Computing Lab, Department of Mathematics, Politecnico di Milano,city=Milan, country=Italy
[…]
[inst2]organization=Division of Pediatric Infectious Diseases, Department for Woman and Child Health, University of Padova,city=Padova, country=Italy
[…]
[inst3]organization=Unit of Biostatistics, Epidemiology and Public Health, Department of Statistics and Quantitative Methods, University of Milano-Bicocca,city=Milan, country=Italy
[…]
[inst4]organization=National Centre for Healthcare Research and Pharmacoepidemiology, Department of Statistics and Quantitative Methods, University of Milano-Bicocca,city=Milan, country=Italy
[…]
[inst5]organization=HDS - Health Data Science Centre, Human Technopole,city=Milan, country=Italy
1 Introduction
Electronic Health Records contain a vast and continuously growing amount of patient information, but a significant portion of this data remains unstructured in formats like clinical notes, reports, discharge letters and referrals. Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity [1, 2]. While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [3, 4], applying these techniques to healthcare data presents additional challenges, such as a specialized lexicon, a high prevalence of ambiguous abbreviations, frequent misspellings, and a lack of standard templates across hospitals [5, 6]. Additionally, most NLP tools are developed for high-resource languages, making them less effective for languages like Italian, where linguistic resources are scarce [7].
A critical piece of information in medical documents is the reason for a patient’s hospital admission or discharge, typically a disease diagnosis or a description of the patient’s symptoms. This information is critical for both cohort selection and epidemiological analysis. Although an international code system for diseases and symptoms, the ICD-CM [8], exists, physicians often prefer to provide a textual description of their diagnoses. Textual descriptions are faster to write, more detailed, and are often used because many medical document templates do not include structured fields for disease codes [9, 10]. Even when present, it has been observed that ICD codes are not completely reliable [11, 12, 13]. Consequently, identifying patients admitted for specific diseases or symptoms often requires manual inspection of these texts, which is labour-intensive and impractical at scale.
In Italy, every hospitalization or Emergency Room (ER) admission generates a discharge letter. Depending on the type of examinations and the hospital, more structured data may also be recorded. A discharge letter should always contain information about the diseases that affect the patient, but the complexity of these documents and the lack of standard templates pose a significant challenge for automatic data extraction. To address this issue, this study aims to develop an NLP pipeline capable of extracting structured information from the unstructured text of Italian discharge letters.
Traditional approaches to this problem typically frame it as a supervised classification task, which requires a substantial amount of annotated data to perform well [14]. However, annotating thousands of discharge letters is time-consuming, especially for clinicians who are already overworked. This highlights a significant gap in the field: the need for methods that can effectively utilize large, unannotated datasets without the burden of manual labelling.
While there have been a few studies on NLP applied to Italian healthcare documents, none have focused specifically on discharge letters. Lanera et al. [15, 16] worked on supervised models to identify cases of varicella, using textual data from specific children’s health project dataset. Similarly, Sciannameo et al. [17] used a BERT-based model to predict rehospitalizations for elderly people from the text of medication prescriptions and hospitalization diagnoses. Hammami et al.  [18] classified pathology reports for cancer morphology, developing a rule-based system. Other studies have concentrated on extracting events and entities, such as Viani et al. [19, 20], who analyzed cardiology reports using ontologies and recurrent neural networks.
Significant research on discharge letters has been done in the English language, often using the MIMIC dataset [21, 22] (see [23, 24, 25, 26] for some examples).
Research has also been conducted on discharge letters in other under-resourced languages, such as Bulgarian [27, 28], Dutch [29] and Swedish [30].
All these studies have relied either on large annotated datasets or on the development of complex, problem-specific rule-based systems. In contrast, our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than traditional rule-based systems.
The main contribution of this paper is the first NLP-based pipeline for the weakly-supervised classification of Italian discharge letters, offering a scalable solution to a long-standing challenge in healthcare data management.
The results presented in this paper pertain to a clinically relevant case study aimed at identifying bronchiolitis, one of the most relevant diseases that can determine hospitalizations in young children [31]. An automated model for this purpose can significantly aid in patient cohort selection for research, reducing the manual workload for physicians. The dataset for this case study comprises 33’176 discharge letters related to children from 44 different hospitals in the Veneto region, one of the most populated Italian regions.
The rest of the paper is structured as follows: in Section 2, we discuss the methodology we developed and the data of our case study; in Section 3, we present the results; in Section 4 we discuss them, together with possible future developments, and in Section 5 we summarize the main findings.
Statement of significance
Problem Identifying diagnoses from discharge letters is relevant for both cohort selection and epidemiological analysis but requires large annotated datasets, rarely available in this domain.
What is Already Known Supervised learning methods for diagnosis identification are well-studied, with state-of-the-art methods based on transformers. Unsupervised methods are mainly rule-based algorithms that lack generalizability.
What this Paper Adds This paper introduces a weakly-supervised NLP pipeline that eliminates the need for manual annotations, offering a scalable solution for diagnosis identification. The approach is validated on a case study with a large dataset of Italian discharge letters and demonstrates strong performance and potential generalizability.
2 Methods
In this section, we describe in detail the novel pipeline we propose for the weakly-supervised diagnosis identification. The pipeline is sketched in Figure 1 and it can be divided into three main steps:
  1. 0
    BPT - Pre-training of a BERT-based model for Italian medical documents.
  2. 1
    SAL - Semi-automatic labelling of the dataset.
  3. 2
    WLC - Classification of discharge letters using the weakly-labelled dataset.
Step 0 (BPT) is described in Section 2.0.1. It aims to build a model able to produce embeddings of Italian medical documents that are exploited by both Step 1 (SAL) and Step 2 (WLC), described in Section 2.0.2 and 2.0.3.
Figure 1: Pipeline schema. Diagram of the full pipeline for weakly-supervised classification of discharge letters
2.0.1 Pre-training of a BERT-based model for Italian medical documents (BPT)
Statistical and machine learning models for textual data require a mathematical representation of the text. There are three main types of possible representations: vocabulary-based, static neural network-based embeddings and contextual neural network-based embeddings. Vocabulary-based representations produce a high dimensional vector whose size is equal to the corpus vocabulary size and whose content is sparse. Their main issue is that they cannot capture the semantic relationships between words, i.e., the fact that different words may have the same or similar meaning, possibly also depending on the context in which they appear. These issues are overcome by neural network-based embeddings. In particular, static word embeddings, such as Word2Vec [32], produce embedding vectors that are permanently fixed for each term. In contrast, contextual embeddings are related to the context in which the word appears, capturing more semantics at the price of requiring an inference through a neural network each time the embedding has to be computed. The state-of-the-art model for contextual embeddings is BERT (Bidirectional Encoder Representations from Transformers) [33], a deep learning model based on the transformer architecture [34]. Different versions of BERT have been developed for different languages and domains. In particular, for the English language, there are various versions of BERT related to the medical domain, such as BioBERT [35] and ClinicalBERT [36].
For the Italian language, the only available versions of BERT are trained on general-domain documents crawled from the web [37]. Thus, we fine-tuned the last Italian version of BERT, Umberto, using the E3C Corpus [38], a multi-lingual dataset composed of documents in the medical domain, particularly clinical narratives from various sources: journal papers, admission tests for specialities in medicine, patient information leaflets for medicines, and abstracts of theses in medical science. The dataset contains documents in five languages; a summary of its composition is provided in Supplementary Material A3. This is the only publicly available dataset that includes Italian medical documents. To utilize the entire dataset, which amounts to 564 MB, we automatically translated the non-Italian documents using Google Translate’s APIs (Step 0.1 in Figure 1). We further pre-trained the Umberto model with this dataset, adopting the usual masked-word prediction task (Step 0.2 in Figure 1). This additional pre-training of the Umberto model over the E3C corpus involved 3 epochs of training (until the validation perplexity did not improve further) with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
2.0.2 Semi-automatic labelling pipeline (SAL)
We aim to develop a classification system for discharge letters that can work on an unlabelled dataset. To do so, we aim to construct weak labels that a supervised classification model can subsequently use. This section presents a methodology to construct weak labels by extracting diagnosis strings from discharge letters and clustering them. This allows the creation of groups of diagnoses related to the same disease or symptoms that can be exploited as labels. This step corresponds to Step 1 (SAL) of Figure 1.
Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on the BERT model.
In the following paragraphs, we describe in detail these sub-steps.
Extraction of diagnosis strings
The first step, corresponding to block 1.1 in Figure 1, is the extraction of diagnosis strings. Italian discharge letters do not follow a standardized template, as each hospital has its own, and even different departments within the same hospital may personalize the template further. Nevertheless, we have observed that these templates often include a designated section for the diagnosis, where the doctor should report the patient’s admission diagnosis. This diagnosis indication should contain, in most cases, the necessary information to determine the label for the discharge summary. However, there are three main challenges in handling this diagnosis indication:
  1. 1.
    Not all discharge summaries include this indication, as it may be absent or left empty by the physician.
  2. 2.
    Different templates use varying terminologies and position this indication differently in the text.
  3. 3.
    Diagnosis strings are typically written as free text, with the same disease described by different strings and different levels of detail
The first point is the primary reason why our pipeline focuses on classifying the text of a discharge letter rather than just the diagnosis string. To address the second point, we employ regular expressions to detect the portion of text where they are present. Regular expressions provide a powerful and flexible approach for pattern matching, making them suitable for capturing a template field with limited variations, such as the diagnosis indication in our dataset. We reported in Supplementary Material C the expressions we defined to extract the diagnosis strings.
After this step, we have, for the majority of the letters, a string, typically a short sentence, explaining the reasons for the patient’s admission. Due to the third point mentioned before, it is not possible to directly exploit them as labels. To do this, it is necessary to cluster them, as detailed in the following paragraph.
Diagnosis Clustering
Once the diagnosis strings are extracted, we use a tailored clustering pipeline to group them meaningfully, aiming at identifying clusters related to the various types of diseases and symptoms. This step corresponds to block 1.2 in Figure 1 and is expanded in Figure 2.
Figure 2: Clustering section of the pipeline. Detailed diagram of the clustering pipeline for diagnosis strings, corresponding to Step 1.2 of the full pipeline depicted in Figure 1.
Punctuation removal, digits removal and abbreviation expansion are applied to the diagnosis strings (step 1.2.1). This pre-processing aims to facilitate the subsequent construction of a text representation that can be representative of the semantics of the text. In particular, punctuation signs and digits are irrelevant for discriminating between diagnoses and can be safely removed. The problem of abbreviations is a common problem in medical documents: a wide variety of acronyms exist in the medical domain, and their meaning can also be context-dependent. We have extracted all the words with two and three letters from our dataset of diagnosis strings and expanded those that we found to be medical abbreviations.
After this pre-processing, the pipeline needs to construct an embedding of the text that a clustering algorithm will subsequently use. This step (1.2.2) exploits the BERT model we discussed in Section 2.0.1 (BPT).
The BERT-based models produce embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
After this step, a clustering algorithm can be applied to the resulting embedding vectors (Step 1.2.4). We selected HDBSCAN since it provides many relevant advantages: it does not require defining a priori the number of clusters, it does not necessarily cluster all the data, and it can identify clusters of varying shapes and sizes.
To effectively group strings related to different diseases and symptoms, the clustering must prioritize the intra-cluster similarity against the inter-cluster dissimilarity. This has the drawback of creating multiple clusters associated with each disease/symptom that are very compact internally, but that should be merged. A second-level clustering can help merge these clusters that are initially separate, simplifying the identification of the weak labels. To this purpose, a suitable representation for clusters needs to be produced. We employ a keyword representation in which, for each cluster, we extract words that appear with significant frequency in the cluster and do not appear with considerable frequency over the entire dataset (Step 1.2.5). These keywords can be represented with BERT embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
Mapping clusters to diseases
Step 1.3 consists of mapping the clusters to the weak labels that will be used by the classification model. This requires analyzing the diagnosis clusters produced by Step 1.2, using their keyword-based representation (Step 1.2.5). The selection can be fully automated with keyword matching or manual. To automate this step, we propose the following procedure:
  1. 1.
    Definitions: identify a set of definitions for the disease or symptoms of interest. Each definition includes one or more positive keywords (directly related to the disease) and, if necessary, negative keywords (terms indicating alternative diagnoses).
  2. 2.
    Cluster selection: for each cluster, select it if it contains all the positive keywords and none of the negative keywords, according to at least one of the definitions.
This procedure requires minimal manual effort upfront for defining keywords and allows for a more precise selection than other methods, such as using similarity scores between embeddings of definitions and cluster representations.
A keyword-based selection at this step is more straightforward and effective than applying keywords directly to the diagnosis strings extracted in Step 1.1. The original strings are often too heterogeneous, leading to false negatives due to varied terminology and false positives from irrelevant information (e.g., negations, past medical issues). Similarly, manually reviewing the original diagnosis strings is impractical due to their large number.
2.0.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
The classification pipeline that exploits the weakly labelled data produced by Step 1 corresponds to Step 2 in Figure 1 (WLC). A pre-processing step (Step 2.1) is necessary to remove headers and footers that contain information that is non-relevant to our analysis. Supplementary Material B reports details about this. The classification model we adopted is based on the fine-tuned version of the Umberto model that we presented in Section 2.0.1 (BPT). In this case, the transformer-based model is enhanced with a linear classification layer that outputs the probability of having the disease. It is trained with the discharge letters’ text and the weak labels assigned to the letters based on the clusters to which their diagnoses have been assigned (Step 2.2).
Given the limitation of BERT-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it). This is not a relevant limitation, considering that 90% of discharge letters in our dataset are within 512 tokens after pre-processing. In the case of datasets with longer texts, it is possible to chunk them in portions no longer than 512 tokens, process them in parallel and combine their outputs, taking the maximum probability.
2.1 Training and evaluation
The BERT-based classifier is trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for six epochs. Initially, only the last classification layer is trained, and after two epochs, the last BERT layer is unfrozen. After four epochs, the last but one BERT layer is unfrozen. These hyperparameters have been selected with grid search.
We cannot perform a precise separate evaluation of all the pipeline steps since we do not have ground truth labels related to diagnosis strings and all their clusters. Because of this, we assess the performance of the whole pipeline on the classification task, implicitly evaluating all the previous steps. For the clustering, we can perform a specific evaluation limited to the clusters related to our bronchiolitis gold labels.
[…]
The classification results are evaluated in terms of precision, recall and F1-score for the positive class and area under the receiver operating characteristic curve (AUC). These metrics are computed for both our weak labels and the gold standard bronchiolitis labels. Evaluating against weak labels allows for an assessment of the classification model’s performance alone, while evaluation against gold standard labels provides a comprehensive view of the pipeline’s effectiveness. We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised classifier using weak labels. We compute the metrics with 10-fold stratified cross-validation.
Considering that our weak labels are derived from clusters of diagnosis strings that we extracted from discharge letters, this might bias the classification models into looking only into the diagnosis sections of the letters. To assess the presence of this bias and to verify whether the models can exhibit good performance with the remaining text of the string, we test the models in two settings:
  1. 1.
    With the original letters, possibly including diagnosis strings
  2. 2.
    Removing the diagnosis strings exploited to assign weak labels
Since our dataset contains data from different hospitals and different LHUs, we verify the presence of differences among them. Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach. For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
We also compare results between the groups of pediatric vs non-pediatric ER/Department admissions. In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
3 Results
In this section, we present the results of our pipeline for the bronchiolitis case study, for the different stages that compose it. In particular, Section 3.1 describes the dataset, Section 3.2 presents results related to the extraction of diagnosis strings from discharge letters, Section 3.3, covers the clustering pipeline and Section 3.4 presents the final results on the identification of bronchiolitis.
3.1 Dataset
The dataset comprises 33’176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020. The dataset covers 44 different hospitals in the region, which can be grouped into 9 Local Health Units (LHUs) and 2 university hospitals. Each discharge letter in the dataset includes the complete text of the letter, along with the date of the letter and anonymized identifiers of the patient. A manual review of discharge letters was conducted to extract relevant diagnoses of bronchiolitis, resulting in 335 positive cases within the dataset. This review was performed by a data manager with a clinical background, with the support of clinicians for uncertain cases. These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training.
We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (e.g., pediatria, pediatrico, pediatrica) in the header of the letters. This marked 10’340 records (31.17%) as related to pediatric ERs/Departments. Supplementary Material A1 reports additional plots with data distributions among hospitals and LHUs. In particular, ten hospitals cover more than 85% of the records.
The median length of the letters is 4515 characters (1st-3rd quartiles 3240-5806). An example of a fictitious letter is reported in Supplementary Material A2.
3.2 Diagnosis extraction
Our diagnosis extraction module (Step 1.1) successfully extracted a diagnosis sentence from 89% of the discharge letters in the dataset. We extracted 15’038 unique diagnosis sentences, of which 75% occur only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms. A few of the extracted strings include the ICD9-CM codes and their standard textual description. Specifically, 737 diagnosis strings (4.9%) contain an ICD9-CM code, covering 3’436 discharge letters (10.4%). The most common strings are the simplest, while more articulate ones have single occurrences. The median length of the extracted diagnoses strings is 49 characters, with 1st and 3rd quartiles of 17 and 156 characters. Among the 335 bronchiolitis-labelled letters, we extracted 138 unique diagnosis strings from 306 letters.
3.3 Diagnosis clustering
In the first clustering level, the 15’038 extracted diagnosis strings were grouped into 590 clusters. Cluster sizes varied significantly, from a large cluster of 270 strings to several smaller clusters containing 5 strings. Table 1 presents the 20 largest clusters, including their sizes and descriptive keywords. The subsequent second-level clustering reduced the number of clusters to 135.
For our case study on bronchiolitis identification, Table 2 details the main clusters related to respiratory infections, including the number of bronchiolitis cases they cover. Four clusters (#1, #2, #3, #5) contain the keyword “bronchiolitis”. Three include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis. These four clusters together cover 65.38% of all bronchiolitis cases. Additionally, 13 clusters include the keyword “bronchospasm”. According to physicians who annotated the data, bronchospasm is associated with bronchiolitis only if accompanied by fever. Only one of these 13 clusters (#4) has both “bronchospasm” and “fever” keywords, and indeed, it contains a high percentage (70%) of bronchiolitis, accounting for an additional 15% of total bronchiolitis cases. Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two definitions of {bronchiolitis} and {bronchospasm, fever}. Among the respiratory infection clusters, three other clusters contain bronchiolitis cases (#6, #7, #8), but they are in a limited number.
For the second-level clustering, Table 3 presents clusters related to respiratory infections. Some clusters from Table 2 have been merged: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections. The selection of weak labels for training remains unchanged since, looking at their keywords, we select second-level clusters #1, #2, #3, #4 that cover the same records as the clusters we would have chosen at the first level. The main advantage of second-level clustering is the substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying weak labelling, especially if manual verification is involved.
This cluster selection for weak labels covers 80% of the original bronchiolitis cases. To understand the reason for the false negatives, we reviewed twenty letters labelled as bronchiolitis in the gold standard but not covered by our weak labels. We found 10 cases where the diagnosis pertains to other symptoms or disease (e.g., otitis, dyspnea, fever), but the bronchiolitis is mentioned in the letter, at least as a suspect condition; 4 cases where bronchiolitis is mentioned in the diagnosis alongside other diseases, leading to assignment to a cluster related to the other disease; 3 cases of respiratory issues with no explicit mention of bronchiolitis, respiratory syncytial virus, or bronchospasm and fever; 2 cases with no diagnosis string present; 1 case where a family member has bronchiolitis, but the patient’s diagnosis is of a different disease. These results suggest that the absence of some gold labels in our weak labels primarily arises from the lack of explicit bronchiolitis indications, with only a few cases that could have been detected but were missed. This highlights the importance of using the full discharge letters for a model aiming to identify diseases accurately. Furthermore, there might be a few errors in the gold labels themselves.
We cannot perform a complete evaluation of the clusters produced using our Umberto-E3C model against those produced using the original Umberto model, but we can compare the distribution of gold bronchiolitis labels on the two sets of clusters. Supplementary Material D reports the clusters obtained using the original Umberto model in our clustering pipeline. The number of clusters related to respiratory infections is higher (34 vs 27), and their keywords appear more heterogeneous. Selecting weak labels by applying the same criterion we used on our clusters, we would cover only 47% of bronchiolitis cases with the clusters from the original Umberto model. Moreover, they are less compact, introducing more false negatives in the weak labels.
Table 1: Keyword-based representation of the twenty largest first-level diagnosis clusters
Cluster size Cluster keywords [ITA] Cluster keywords [ENG]
270 flogosi vie phlogosis airways
145 trauma cranico head injury
143 contusione dito mano contusion finger hand
116 gastroenterite verosimile probable gastroenteritis
114 episodio febbre fever episode
94 vie infezione tract infection
93 pronazione dolorosa painful pronation
78 ferita labbro lip wound
75 otite media acuta acute otitis media
70 ferita lacero-contusa lacerocontused wound
66 trauma contusivo blunt trauma
66 broncospasmo flogosi vie bronchospasm pathway inflammation
64 puntura insetto insect bite
62 ferita lacero-contusa lacerocontused wound
59 trauma dx right-handed trauma
58 ustione grado burn degree
54 frattura ferita wound fracture
54 ferita lacero contusa destra right-handed lacerocontused wound
53 insorgenza recente febbre recent onset fever
53 contusione piede destro dito right foot toe contusion
Table 2: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
Cluster # Cluster keywords [ITA] Cluster keywords [ENG]
Within cluster
bronch. %
% of total bronch.
covered
1 bronchiolite lieve mild bronchiolitis 85.19% 6,87%
2 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.14% 44,48%
3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05% 11,94%
4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70,00% 15,02%
5 alimentazione bronchiolite difficolta’ feeding difficulties bronchiolitis 50.00% 2,09%
6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7% 2,39%
7 broncospasmo infezione respiratoria bronchospasm respiratory infection 6.67% 0,30%
8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17% 0,60%
9 flogosi alte vie respiratorie febbre high airways flogosis fever $0$,00%
10 broncospasmo infezione vie respiratorie bronchospasm airways infection $0$,00%
11 acuto broncospasmo acute bronchospasm $0$,00%
12 broncospasmo flogosi vie bronchospasm airways flogosis $0$,00%
13 broncospasmo polmonite bronchospasm pneumonia $0$,00%
14 broncospasmo corso ongoing bronchospasm $0$,00%
15 broncopolmonite sinistra left bronchopneumonia $0$,00%
16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis $0$,00%
17 broncospasmo episodio bronchospasm episode $0$,00%
18 infezione alte vie respiratorie high airways infection $0$,00%
19 flogosi infezione alte vie respiratorie flogosis high airways infection $0$,00%
20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis $0$,00%
21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm $0$,00%
22 broncospasmo otite bronchospasm otitis $0$,00%
23 broncospasmo parainfettivo bronchospasm parainfective $0$,00%
24 broncospasmo virosi bronchospasm virosis $0$,00%
25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory $0$,00%
26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient $0$,00%
27 altre infezioni acute vie respiratorie other acute respiratory infection airways $0$,00%
Table 3: Keyword-based representation of the second-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG] Within cluster bronch. % % of total bronch. covered 1 bronchiolite lieve mild bronchiolitis 1 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.16 % 51.34 % 2 bronchiolite lieve mild bronchiolitis 2 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 % 11.94 % 3 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 % 15.02 % 4 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 5 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 50.00 % 2.09 % 5 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7 % 2.39 % 6 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo infezione vie respiratorie bronchospasm infection airways 2.9 % 0.30 % 10 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo disidratazione lieve bronchospasm mild dehydration 8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 % 0.60 % 8 broncospasmo bronchospasm 11 acuto broncospasmo acute bronchospasm 0.00 % 0.00 % 12 broncospasmo flogosi vie bronchospasm airways flogosis 13 broncospasmo polmonite bronchospasm pneumonia 14 broncospasmo corso ongoing bronchospasm 22 broncospasmo otite bronchospasm otitis 23 broncospasmo parainfettivo bronchospasm parainfective 24 broncospasmo virosi bronchospasm virosis 17 broncospasmo episodio bronchospasm episode 9 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 % 0.00 % 10 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 % 0.00 % 11 infezione vie respiratorie airways infection 18 infezione alte vie respiratorie high airways infection 0.00 % 0.00 % 19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 % 0.00 % 27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 % 0.00 % 26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 % 0.00 % 16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 % 0.00 % 12 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 % 0.00 % 13 alte febbre flogosi respiratorie vie high fever flogosis airways 9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 % 0.00 % 14 broncopolmonite sinistra left bronchopneumonia 15 broncopolmonite sinistra left bronchopneumonia 0.00 % 0.00 %
3.4 Disease classification
[…]
When comparing the classification results on weak labels (W) and gold labels (G), we observe a reduction in precision, affecting both the F1-Score and the AUC. However, recall is higher on gold labels, at least for BERT-based models, whereas simpler models also showed a decrease in recall.
The BERT-based models show minimal differences between settings with and without diagnosis string, indicating that the information needed to identify the diagnosis of bronchiolitis is present in the entire letter text, and these models can effectively recognize it.
As expected, the fully supervised model demonstrated better overall performance, though differences are modest, particularly in recall, which is lower on average than our best weakly-supervised model.
In Table 5, we assess the differences in the results of our best model when selecting different subsets of clusters for the weak labels. In particular, we exclude one of the 1st level clusters we selected at a time. Excluding clusters #5 or #1 does not affect the results, which is reasonable considering their smaller size. Excluding cluster #2 determines a drop in the performance and an increase in the variance, which is expected since it accounts for almost half of the bronchiolitis cases. Removing clusters #3 or #4 has a negative effect, but limited. These results confirm the solidity of our weak labels and the robustness of the model with respect to small changes in them.
Table 6 compares metrics at the hospital level, revealing notable differences. Some hospitals, such as H-0 and H-3, are more represented in the dataset, which might explain their superior performance. However, hospitals with a similar number of records show considerable performance variability, suggesting that reporting practices in discharge letters might influence disease detection. A comparison between records from pediatric vs non-pediatric ER/Departments revealed significant differences (gold AUC $80.89\pm 6.37$ vs $71.07\pm 4.57$, see Supplementary Material E for details), but there are also hospitals with pediatric ER and low performances, such as H-4. Similar findings are reported in Supplementary Material E for the different LHUs.
Table 4: 10-fold CV results for classification, comparing with rule-based baselines, fully supervised approaches and other classification modules in our pipeline. Std. Dev. in parenthesis. G = gold labels, W = weak labels.
With diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 82.30 (8.94) 75.46 (6.25) 77.33 (4.06) 82.30 (4.99) 72.75 (7.50) 84.59 (6.48) 75.06 (3.72) 77.68 (5.73)
Umberto 80.67 (5.99) 73.48 (7.12) 76.91 (5.97) 82.56 (4.79) 62.29 (6.24) 75.99 (4.69) 68.46 (2.77) 69.82 (4.54)
LSTM 68.57 (6.48) 64.97 (11.96) 65.54 (5.66) 67.50 (3.99) 57.39 (11.37) 60.86 (13.67) 55.08 (8.07) 63.89 (5.28)
Rule-based full text 70.18 87.05 77.70 - 65.57 68.76 67.13 -
Rule-based diagnosis 73.88 86.38 79.64 - 74.61 67.42 70.83 -
Umberto-E3C Supervised - - - - 80.30 (7.23) 81.52 (6.70) 80.91 (2.81) 86.45 (2.82)
Without diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 84.34 (5.19) 73.04 (10.53) 75.15 (4.27) 84.91 (3.50) 70.15 (6.40) 79.00 (5.82) 74.28 (4.10) 77.56 (3.11)
Umberto 80.01 (6.84) 68.68 (10.55) 73.90 (5.36) 80.81 (6.24) 61.41 (6.09) 76.69 (4.22) 68.18 (3.62) 70.49 (3.89)
LSTM 64.98 (6.89) 56.21 (18.88) 58.80 (13.52) 63.00 (9.24) 50.14 (8.31) 63.97 (13.52) 53.03 (9.69) 59.02 (7.36)
Rule-based full text 65.54 79.16 71.69 - 62.33 60.05 61.10 -
Umberto-E3C Supervised - - - - 75.56 (5.08) 82.55 (9.14) 78.90 (4.54) 84.05 (4.26)
Table 5: 10-fold CV results for classification, considering different groups of 1st level clusters for weak labels. Std. Dev. in parenthesis. G = gold labels, W = weak labels
Clusters for weak labels
(1st level)
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
#1. #2. #3. #4 78.39 (7.04) 73.35 (6.21) 75.79 (4.65) 83.02 (4.16) 70.75 (6.48) 81.27 (6.99) 75.65 (4.29) 75.49 (5.57)
#1. #2. #4. #5 73.69 (8.89) 83.20 (9.16) 78.16 (7.80) 80.71 (7.01) 68.42 (8.70) 75.09 (8.58) 71.60 (7.26) 72.67 (6.87)
#2. #3. #4. #5 83.49 (7.87) 74.66 (6.88) 78.82 (4.26) 84.02 (4.43) 68.54 (6.30) 82.35 (4.94) 74.81 (3.82) 75.83 (5.22)
#1. #3. #4. #5 93.73 (6.85) 78.11 (8.47) 85.19 (7.06) 90.91 (7.35) 57.29 (12.97) 59.03 (15.94) 58.15 (12.45) 68.90 (10.04)
#1. #2. #3. #5 80.45 (8.90) 74.62 (8.57) 77.43 (5.76) 85.95 (5.29) 67.86 (6.36) 73.42 (8.64) 70.53 (5.33) 70.22 (3.10)
Table 6: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
Hospital Dataset %
Pediatric ER
Bronchiolitis
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
H-0 31.41 87.30 0.76 77.46 78.57 78.01 90.40 81.40 63.64 71.43 86.56
H-1 14.56 0.00 1.08 63.64 71.79 79.43 85.32 66.67 66.67 66.67 77.95
H-2 11.63 1.00 0.78 80.00 72.73 76.19 92.16 50.00 71.43 58.82 83.98
H-3 10.23 12.00 1.71 88.68 88.68 88.68 89.45 81.40 81.40 81.40 88.12
H-4 7.43 22.00 1.22 92.31 35.29 51.06 64.32 84.21 37.21 51.61 57.34
H-6 2.59 9.40 2.09 100.00 31.25 47.62 65.63 100.00 33.33 50.00 70.48
4 Discussion
This study introduces a novel pipeline that shows how weakly-supervised learning can be effectively conducted in the clinical natural language processing domain. This is the first study to address the problem of diagnosis identification from discharge letters in the Italian language, and while we validated the method using a case study on bronchiolitis, the pipeline is designed to be applicable to any disease. Our case study shows excellent results, outperforming existing unsupervised approaches, and underscores the potential for broader application in clinical settings.
We observed minimal performance differences between the metrics computed on weak and gold labels (1-2 % of F1-Score, 5-7 % of AUC), confirming that the first step of our pipeline (SAL) effectively identifies relevant cases. Despite the inevitable noise that the weak labels introduce, this is limited enough to allow the model to perform well when evaluated against gold labels. Additionally, the variations between settings where we retained the diagnosis strings and those where we removed them are negligible. This indicates that the model is effectively learning from the entire text of the discharge letters, making it particularly advantageous in situations where automatic extraction of diagnosis strings is not possible. Our manual exploration of some weak labels further revealed that diagnosis strings do not always mention the disease responsible for hospitalization. In some cases, they only refer to symptoms, and in cases of multiple diseases, not all may be reported. These findings highlight the necessity for a classification model based on the complete text of discharge letters, as the one we developed.
In contrast, rule-based approaches exhibited poor performance, mainly due to the fact that diseases like bronchiolitis are not always explicitly mentioned by name in the text. Even when mentioned, they might refer to discarded hypotheses or contacts with affected individuals, leading to false positives. While a more sophisticated rule-based system could address some of these issues, such approaches lack the scalability and generalizability of our pipeline. Conversely, our pipeline can be easily adapted to other diseases, which makes it a versatile tool for clinical NLP.
Our fine-tuned version of the Umberto model showed the best results both in the classification and the clustering modules, demonstrating a better understanding of the Italian medical domain when compared with the general-domain version.
Compared to a fully supervised method, our weakly-supervised pipeline maintained competitive performance, with only a limited and acceptable loss ($\sim 5\%$ of F1-Score, $\sim 6\%$ of AUC). This is particularly notable given the significant time and effort required for manual annotation in fully supervised setups. The pipeline has also shown robustness to the cluster selection for the weak labels. These results underscore the potential of weakly-supervised approaches in clinical natural language processing, a domain where the scarcity of labelled data is the most relevant challenge.
While our results are promising, we must acknowledge some limitations in our study that could be addressed in future work. The absence of a gold standard in labelling is a common challenge in clinical NLP, which underscores the value of weakly-supervised methods like ours. Our gold labels for bronchiolitis were manually assigned by individual physicians, making it impossible to verify inter-annotator agreement and increasing the risk of errors.
The multi-centric nature of our dataset, though a strength in its diversity, also presents an imbalance in the number of records from different hospitals and Local Health Units. Our analysis highlighted relevant performance differences among hospitals and LHUs, warranting further investigation. Testing the pipeline on a more balanced dataset would strengthen conclusions about these performance differences. Moreover, this could facilitate the development of a hierarchical transformer-based classification model that considers hospital/LHU context in the input, potentially improving model performance.
We also observed relevant differences between pediatric vs. non-pediatric ERs/Departments. Regarding this, we need to consider that pediatric admissions were identified using specific keywords. This method could miss some pediatric cases, although such instances should be limited. Future work could use data that directly includes this information as a structured field, ensuring its accuracy and completeness.
Although our case study focused on bronchiolitis, the diagnostic clusters we identified appear capable of recognizing a broad range of diseases. Validating the pipeline on other diseases would further demonstrate its robustness and generalizability. Developing a multi-disease model, particularly for conditions that share similar symptoms, could provide broader applicability and enhance the model’s utility. Tailored rule-based models often fail in such scenarios, making a machine learning-based approach more effective.
Another important direction for future research is enhancing the interpretability of our model. Identifying the portions of the discharge letter that are most relevant to the model’s output would help clinicians understand and trust the model’s decisions. Techniques related to the area of explainable AI could be employed to enhance interpretability, ensuring the model’s reasoning process is transparent and understandable. This would also help mitigate biases that could potentially be present in the training data, further enhancing the model’s reliability.
5 Conclusions
We have developed and validated the first weakly-supervised pipeline to analyze discharge letters in the Italian language, with a case study focusing on bronchiolitis. The pipeline achieved excellent results, outperforming rule-based approaches. Our fine-tuned BERT-based model showed superior performance in both the clustering and the classification module compared to a general-domain BERT model. We also assessed significant differences between hospitals/LHUs and pediatric vs. non-pediatric ERs, which suggest further investigations.
However, while the case study showcased the method’s potential, the pipeline itself is designed to be adaptable to any disease, making it a valuable tool for researchers interested in extracting diagnoses from unstructured data and for epidemiological monitoring of diseases.
Despite a small performance gap compared to fully supervised approaches, the weakly-supervised pipeline addresses real-world constraints, particularly the scarcity of labelled data in clinical NLP. Transformer-based models showed strong performance even without gold-labelled data, underscoring the viability of weakly-supervised methods in overcoming data limitations in the medical domain.
Declaration of competing interests
The authors declare that they have no competing interests.
Data statement
Data used in this study are collected and accessed through Pedianet, an Italian network of paediatricians. Data anonymization is performed in accordance with Italian regulations. Each anonymized record is assigned a unique numerical identifier. The data is then transmitted monthly to a centralized database located in Padua, Italy, for validation purposes. Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide consent for their children’s anonymized data to be used for research purposes. Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te. Srl, the legal owner of Pedianet.
Data used in this study are not publicly available due to their confidential nature. The code developed for their analysis is available by the authors upon request.
CRediT authorship contribution statement
Vittorio Torri: Methodology, Software, Investigation, Writing - Original Draft Elisa Barbieri: Conceptualization, Data Curation, Writing - Review & Editing Anna Cantarutti: Conceptualization, Data Curation, Writing - Review & Editing Carlo Giaquinto: Conceptualization, Supervision Francesca Ieva: Conceptualization, Methodology, Writing - Review & Editing, Supervision
Acknowledgments
The present research has been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027.
Elisa Barbieri declares that her position is funded within the project “INF-ACT - One Health Basic and translational Research Actions addressing Unmet Needs on Emerging Infectious Diseases”, ID MUR PE_00000007, under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2 – CUP C93C22005170007.
The authors thank all the family paediatricians collaborating in Pedianet: Eva Alfieri, Michela Alfiero Bordigato, Angelo Alongi, Biagio Amoroso, Rosaria Ancarola, Barbara Andreola, Giampaolo Anese, Roberta Angelini, Maria Grazia Apostolo, Giovanna Argo, Giovanni Avarello, Lucia Azzoni, Maria Carolina Barbazza, Patrizia Barbieri, Gabriele Belluzzi, Eleonora Benetti, Filippo Biasci, Franca Boe, Stefano Bollettini, Francesco Bonaiuto, Anna Maria Bontempelli, Matteo Bonza, Sara Bozzetto, Andrea Bruna, Ivana Brusaterra, Massimo Caccini, Laura Calì, Sonia Camposilvan, Laura Cantalupi, Luigi Cantarutti, Chiara Cardarelli, Giovanna Carli, Sylvia Carnazza, Massimo Castaldo, Stefano Castelli, Monica Cavedagni, Giuseppe Egidio Cera, Chiara Chillemi, Francesca Cichello, Giuseppe Cicione, Carla Ciscato, Mariangela Clerici Schoeller, Samuele Cocchiola, Giuseppe Collacciani, Valeria Conte, Roberta Corro’, Rosaria Costagliola, Nicola Costanzo, Sandra Cozzani, Giancarlo Cuboni, Giorgia Curia, Caterina D’alia, Vito Francesco D’Amanti, Antonio D’Avino, Roberto De Clara, Lorenzo De Giovanni, Annamaria De Marchi, Gigliola Del Ponte, Tiziana Di Giampietro, Giuseppe Di Mauro, Giuseppe Di Santo, Piero Di Saverio, Mattea Dieli, Marco Dolci, Mattia Doria, Dania El Mazloum, Maria Carmen Fadda, Pietro Falco, Mario Fama, Marco Faraci, Maria Immacolata Farina, Tania Favilli, Mariagrazia Federico, Michele Felice, Maurizio Ferraiuolo, Michele Ferretti, Mauro Gabriele Ferretti, Paolo Forcina, Patrizia Foti, Luisa Freo, Ezio Frison, Fabrizio Fusco, Giovanni Gallo, Roberto Gallo, Andrea Galvagno, Alberta Gentili, Pierfrancesco Gentilucci, Giuliana Giampaolo, Francesco Gianfredi, Isabella Giuseppin, Laura Gnesi, Costantino Gobbi, Renza Granzon, Mauro Grelloni, Mirco Grugnetti, Antonina Isca, Urania Elisabetta Lagrasta, Maria Rosaria Letta, Giuseppe Lietti, Cinzia Lista, Ricciardo Lucantonio, Francesco Luise, Enrico Marano, Francesca Marine, Lorenzo Mariniello, Gabriella Marostica, Sergio Masotti, Stefano Meneghetti, Massimo Milani, Stella Vittoria Milone, Donatella Moggia, Angela Maria Monteleone, Pierangela Mussinu, Anna Naccari, Immacolata Naso, Flavia Nicoloso, Cristina Novarini, Laura Maria Olimpi, Riccardo Ongaro, Maria Maddalena Palma, Angela Pasinato, Andrea Passarella, Pasquale Pazzola, Monica Perin, Danilo Perri, Silvana Rosa Pescosolido, Giovanni Petrazzuoli, Giuseppe Petrotto, Patrizia Picco, Ambrogina Pirola, Lorena Pisanello, Daniele Pittarello, Elena Porro, Antonino Puma, Maria Paola Puocci, Andrea Righetti, Rosaria Rizzari, Cristiano Rosafio, Paolo Rosas, Bruno Ruffato, Lucia Ruggieri, Annamaria Ruscitti, Annarita Russo, Pietro Salamone, Daniela Sambugaro, Luigi Saretta, Vittoria Sarno, Valentina Savio, Nico Maria Sciolla, Rossella Semenzato, Paolo Senesi, Carla Silvan, Giorgia Soldà, Valter Spanevello, Sabrina Spedale, Francesco Speranza, Sara Stefani, Francesco Storelli, Paolo Tambaro, Giacomo Toffol, Gabriele Tonelli, Silvia Tulone, Angelo Giuseppe Tummarello, Cristina Vallongo, Sergio Venditti, Maria Grazia Vitale, Concetta Volpe, Francescopaolo Volpe, Aldo Vozzi, Giulia Zanon, Maria Luisa Zuccolo.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT to improve the readability and clarity of some sentences. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.